OpenStudy (anonymous):

Multiply. Simplify if possible. Is my answer to the following problem correct? (sqrt7 -sqrt8)(sqrt7 +sqrt8) is the answer -14+sqrt(8)8?

OpenStudy (anonymous):

\[(\sqrt7 -\sqrt8)(\sqrt7 +\sqrt8)\]

OpenStudy (anonymous):

7-8=-1 is the answer

OpenStudy (anonymous):

wow! I was waaaay off. lol

OpenStudy (anonymous):

you are using \[(a+b)(a-b)=a^2-b^2\]

jimthompson5910 (jim_thompson5910):

Use the idea that \[(A-B)(A+B)=A^2-B^2\] to get \[(\sqrt{7}-\sqrt{8})(\sqrt{7}+\sqrt{8})=(\sqrt{7})^2-(\sqrt{8})^2=7-8=-1\]

OpenStudy (anonymous):

Just multiply them or use the formula \[a ^{2}-b ^{2}=(a+b)(a-b)\]

OpenStudy (anonymous):

so the answer is -1 or 7? I'm confused again lol

OpenStudy (anonymous):

-1 never mind :)

OpenStudy (anonymous):

so if I did \[(\sqrt2)(-2\sqrt2 +3)\] the answer is not 3sqrt 2-4 is it?

jimthompson5910 (jim_thompson5910):

\[\sqrt{2}(-2\sqrt{2}+3)\] \[-2\sqrt{2}*\sqrt{2}+3\sqrt{2}\] \[-2\sqrt{2*2}+3\sqrt{2}\] \[-2\sqrt{4}+3\sqrt{2}\] \[-2*2+3\sqrt{2}\] \[-4+3\sqrt{2}\] \[3\sqrt{2}-4\] So you got it.

OpenStudy (anonymous):

awesome!